# Evaluating Meta-Regression Techniques: A Simulation Study on Heterogeneity in Location and Time

### v1

@@ 7.2 Future Study
Figure 4: Sample Power Curves: These illustrate the difference of when there are one versus five covariates. Both show Case 2 & $n=100$. The x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 5: Sample Residual MSE Distributions: These illustrate the difference between when there is one versus five covariates. Both show Case 2 & $n=100$, see Figure 16 in Appendix C for the MSE of the covariate $x1$.
Figure 6: Sample Trend Estimates: These illustrate the difference between when there is one versus five covariates for Case 2 & $n=100$. Both show the true trend value as a vertical line.
Small versus Large Time Effect
Figure 7: Sample Power Curves: These illustrate a small versus large time effect (respectively Case 2 on the left and Case 8 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 8: Sample Residual MSE Distributions: These illustrate a small versus large time effect (respectively Case 2 on the left and Case 8 on the right). Both show $n=100$ for 1 covariate, see Figure 17 in Appendix C for the MSE of the covariate.
Figure 9: Sample Trend Estimates: These illustrate the difference between when there is a small versus large time effect (respectively corresponding to Case 2 on the left and Case 8 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$
Small versus Large Location Effect
Figure 10: Sample Power Curves: These illustrate a small versus large time effect (respectively Case 1 on the left and Case 2 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 11: Sample Residual MSE Distributions: These illustrate a small versus large time effect (respectively Case 1 on the left and Case 2 on the right). Both show $n=100$ for 1 covariate, see Figure 18 in Appendix C for the MSE of the covariate.
Figure 12: Sample Trend Estimates: These illustrate the difference between when there is a small versus large time effect (respectively corresponding to Case 1 on the left and Case 2 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$
@@
Figure 13: Sample Power Curves: These illustrate less versus more locations (respectively corresponding to Case 2 on the left and Case 6 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 14: Sample Residual MSE Distributions: These illustrate less versus more locations (respectively corresponding to Case 2 on the left and Case 6 on the right). Both show $n=100$ for 1 covariate, see Figure 19 in Appendix C for the MSE of the covariate.
Figure 15: Sample Trend Estimates: These illustrate less versus more locations (respectively corresponding to Case 2 on the left and Case 6 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$.
Appendix D - Supporting Performance Criteria
Figure 16: These are the supporting performance criteria for Figure 5, MSE of the residual. No substantial differences are seen here, so the performance criteria used for assessment is in Figure 5 instead.
Figure 17: These are the supporting performance criteria for Figure 8, MSE of the residual. No substantial differences are seen here, so the performance criteria used for assessment is in Figure 8 instead.
Figure 18: These are the supporting performance criteria for Figure 11, MSE of the residual. No substantial differences are seen here, so the performance criteria used for assessment is in Figure 11 instead.
Figure 19: These are the supporting performance criteria for Figure 14, MSE of the residual. No substantial differences are seen here, so the performance criteria used for assessment is in Figure 14 instead.
@@
Appendix G - Detailed Standard Meta-Analysis Methodology
@@
A fixed effects model allows a regression to account for individual-specific effects (also known as cross-section level effects) such as each study, $s$, or location, $l$ (Cameron & Trivedi, 2005; Wooldridge, 2010). If there are individual-specific effects, but fixed effects are not used, then the model faces the key issue of an omitted variable bias. One way fixed effects can be implemented is with a standard regression that models each cross-section as a dummy variable, as long as one of the cross-sections is removed to avoid the dummy variable trap (Wooldridge, 2010). Another way that fixed effects can be implemented is by averaging at the study-level. The fixed effects specification would be most appropriate when inferences are only made within each grouping (not between groupings) and the random effects assumption does not hold. These constraints and underlying assumptions will be reviewed later in this section, but first let us review the framework of a fixed effects model. A basic meta-analysis using fixed effects uses the following framework in Equation 22 below.
@@
$SE(\hat{\beta})=\sqrt{\frac{1}{\sum^{S}_{s=1}\frac{1}{\sigma^{2}_{s}}}}$ (26)
To calculate the above standard error estimate $SE(\hat{\beta})$, $\sigma^{2}_{s}$ denotes the study-level population variance (Domínguez Islas & Rice, 2018).
@@
In contrast, what if individual-specific effects are completely uncorrelated with each of the explanatory variables? This is unlikely in practice, but when the random effects assumption holds, a random effects model should be used instead of a fixed effects model (Wooldridge, 2010; Mundlak, 1978; Hausman & Taylor, 1981; Baltagi, 2005; Clark & Linzer, 2015). Empirically the unlikelihood of the random effects assumption holding has been proven in a variety of ways. One source illustrates this with wage studies to show unobserved characteristics (where data is unavailable like natural ability) might influence wages but are completely uncorrelated with any explanatory variables like experience or education (Wooldridge, 2010). Another source shows how the random effects assumption does not hold in majority of empirical studies in a variety of economic fields including labor, health and education (Baltagi, 2005). This has also been shown in a simulation study which generated simulation data that simulated both short and long panel data typical of social science data (Clark & Linzer, 2015). These constraints and underlying assumptions will be reviewed later in this section, but first let us review the framework of a random effects model.

### v2

@@ 8 Conclusion
Figure 4: Sample Power Curves: These illustrate the difference of when there are one versus five covariates. Both show Case 1 & $n=100$. The x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 5: Sample Trend Estimates: These illustrate the difference between when there is one versus five covariates for Case 1 & $n=100$. Both show the true trend value as a vertical line.
Figure 6: Sample X1 Bias & Variance Distributions: The estimators bias and variance and MSE are respectively shown left to right. This is for when there is one covariate and can be compared with the figure below which is for when there are five covariates. Both show Case 1 & $n=100$.
Figure 7: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there is five covariates and can be compared with the figure above which is for when there is one covariate. Both show Case 1 & $n=100$.
Smaller versus Larger Time Effect
Figure 8: Sample Power Curves: These illustrate a small versus large time effect (respectively Case 1 on the left and Case 7 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 9: Sample Trend Estimates: These illustrate the difference between when there is a small versus large time effect (respectively corresponding to Case 1 on the left and Case 7 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$
Figure 10: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there is a small time effect (Case 1) and can be compared with the figure below which is for a large time effect (Case 7). Both show $n=100$ for 1 covariate.
Figure 11: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there is a large time effect (Case 7) and can be compared with the figure above which is for a small time effect (Case 1). Both show $n=100$ for 1 covariate.
Smaller versus Larger Location Effect
Figure 12: Sample Power Curves: These illustrate a small versus large location effect (respectively Case 1 on the left and Case 2 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 13: Sample Trend Estimates: These illustrate the difference between when there is a small versus large location effect (respectively corresponding to Case 1 on the left and Case 2 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$
Figure 14: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there is a small location effect (Case 1) and can be compared with the figure below which is for a large location effect (Case 2). Both show $n=100$ for 1 covariate.
Figure 15: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there is a large location effect (Case 2) and can be compared with the figure above which is for a small location effect (Case 1). Both show $n=100$ for 1 covariate.
@@
Figure 16: Sample Power Curves: These illustrate less versus more locations (respectively corresponding to Case 1 on the left and Case 5 on the right). Both show $n=100$ for 1 covariate, the x-axis shows the power to detect a range of discrepancies from the null hypothesis.
Figure 17: Sample Trend Estimates: These illustrate less versus more locations (respectively corresponding to Case 1 on the left and Case 5 on the right). Both show the true trend value as a vertical line for 1 covariate when $n=100$.
Figure 18: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there are less locations (Case 1) and can be compared with the figure below which is for when there are more locations (Case 5). Both show $n=100$ for 1 covariate.
Figure 19: Sample X1 Bias & Variance Distributions: The estimators bias and variance are respectively shown left to right. This is for when there are more locations (Case 5) and can be compared with the figure above which is for when there are less locations (Case 1). Both show $n=100$ for 1 covariate.
@@
Appendix F - Detailed Standard Meta-Analysis Methodology
@@
A fixed effects model allows a regression to account for individual-specific effects (also known as cross-section level effects) such as each study, $s$, or location, $l$ (Wooldridge, 2010; Cameron & Trivedi, 2005; Baltagi, 2005). If there are individual-specific effects, but fixed effects are not used, then the model faces the key issue of an omitted variable bias. One way fixed effects can be implemented is with a standard regression that models each cross-section as a dummy variable, as long as one of the cross-sections is removed to avoid the dummy variable trap (Wooldridge, 2010; Cameron & Trivedi, 2005; Baltagi, 2005). Another way that fixed effects can be implemented is by averaging at the study-level. The fixed effects specification would be most appropriate when inferences are only made within each grouping (not between groupings) and the random effects assumption does not hold. These constraints and underlying assumptions will be reviewed later in this section, but first let us review the framework of a fixed effects model. A basic meta-analysis using fixed effects uses the following framework in Equation 13 below.
@@
$SE(\hat{\beta})=\sqrt{\left[\sum^{S}_{s=1}\frac{1}{\sigma^{2}_{s}}x_{s}^{\top}x_{s}\right]^{-1}_{jj}}$ (17)
To calculate the above standard error estimate $SE(\hat{\beta})$, $\sigma^{2}_{s}$ denotes the study-level population variance (Domínguez Islas & Rice, 2018). The subscript $jj$ refers to the $j-th$ diagonal elements of the variance-covariance matrix of the estimator, this syntax is used when there are multiple covariates.
@@
In contrast, what if individual-specific effects are completely uncorrelated with each of the explanatory variables? This is unlikely in practice, but when the random effects assumption holds, a random effects model should be used instead of a fixed effects model (Wooldridge, 2010; Clark & Linzer, 2015; Baltagi, 2005; Hausman & Taylor, 1981; Mundlak, 1978). Empirically the unlikelihood of the random effects assumption holding has been proven in a variety of ways. One source illustrates this with wage studies to show unobserved characteristics (where data is unavailable like natural ability) might influence wages but are completely uncorrelated with any explanatory variables like experience or education (Wooldridge, 2010). Another source shows how the random effects assumption does not hold in majority of empirical studies in a variety of economic fields including labor, health and education (Baltagi, 2005). This has also been shown in a simulation study which generated simulation data that simulated both short and long panel data typical of social science data (Clark & Linzer, 2015). These constraints and underlying assumptions will be reviewed later in this section, but first let us review the framework of a random effects model.
@@
Appendix G - Detailed Standard Meta-Analysis Methodology
Location-level Effects: Fixed, Random, and Mixed Effects
Fixed Effects
Equation 23 below represents location-fixed effects.
For location level fixed-effects, each location is indexed by $l=1,...,L$, $y_{l}$ is a $(l\times 1)$ vector for the variable of interest, $\mathbf{x}_{l}^{\top}=(1,x_{l1},x_{l2},...,,x_{lp})$ represents a $(l\times p)$ vector of covariates per location $l$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ represents a $(p\times 1)$ vector of effect sizes (i.e. regression coefficients), and $\varepsilon_{l}$ is a $(l\times 1)$ vector for the error term. Note that $\boldsymbol{\alpha}_{l}$ denotes the fixed effect where $\boldsymbol{\alpha}_{l,FE}=\gamma_{l}D_{l}$ such that $D_{l}$ is a vector of dummy variables per location and $\gamma_{l}$ is a vector of fixed effects for location $l$ (which represents the difference in effect size between each location $l$). The parameter is estimated using the weights of each group (in this case, each location) as seen in Equation 24 below.
$SE(\hat{\beta})=\sqrt{\left[\sum^{L}_{l=1}\frac{1}{\sigma^{2}_{l}}x_{l}^{\top}x_{l}\right]^{-1}_{jj}}$ (25)
To calculate the above standard error estimate $SE(\hat{\beta})$, $\sigma^{2}_{l}$ denotes the location-level population variance (Domínguez Islas & Rice, 2018). The subscript $jj$ refers to the $j-th$ diagonal elements of the variance-covariance matrix of the estimator, this syntax is used when there are multiple covariates.
To estimate the standard error above $SE(\hat{\beta})$, the location-level population variance $\sigma^{2}_{l}$ is needed (Domínguez Islas & Rice, 2018). Grouping by location, $l$, instead of by study, $s$, is a key difference since variation in locations is now explicitly controlled for, instead of being implicitly controlled for by the grouping of studies. The subscript $jj$ refers to the $j-th$ diagonal elements of the variance-covariance matrix of the estimator, this syntax is used when there are multiple covariates.
Random Effects
The simulation will also test location-level random effects, which is similar, but has some slight differences. This also uses Equation 23, but the fixed effect is replaced with the random effect $\boldsymbol{\alpha}_{l,RE}$. This random effect is drawn from $N(0,\sigma_{\alpha}^{2})$ and the error term $\varepsilon_{l}$ is $N(0,\sigma_{\varepsilon}^{2})$ distributed. The estimation of parameters and their standard errors is the same as it was for study-level random effects in Appendix F, just the subscripts change from study-level, $s$ to location-level, $l$ as seen below in Equation 26 & 27.
Mixed Effects
In the equation above, $y_{l}$ denotes a $(l\times 1)$ vector for the dependent variable, $\mathbf{x}_{l}^{\top}=(1,x_{l1},x_{l2},...,,x_{lp})$ representing a $(l\times p)$ vector of covariates per location $l$, $\boldsymbol{\beta}=(\beta_{0},\beta_{1},...,\beta_{p})$ representing a $(p\times 1)$ vector of regression coefficients (i.e. effect sizes), $\boldsymbol{\alpha}_{l,FE}$ denotes the fixed effects vector, $\boldsymbol{\alpha}_{l,RE}$ denotes the random effects vector, and $\varepsilon_{l}$ denotes the $(l\times 1)$ error term vector (Viechtbauer, 2007). Estimating the parameters and their standard errors will be the same as it is for location-level random effects for the same reasons articulated for study-level mixed effects in Appendix F (please see Equation 26 & 27). This is because a mixed effect model is comprised of both random effects as well as fixed effects, and fixed effects provide the foundation to calculate random effects.
Time-level Fixed Effects
Estimation of parameters and their standard errors are the same as it was in Section 4.1 and for study-level fixed effects in Appendix F. The parameter is estimated using the weight of each group (in this case each time) as seen in Equation 29 below.
$SE(\hat{\beta})=\sqrt{\left[\sum^{T}_{t=1}\frac{1}{\sigma^{2}_{t}}x_{t}^{\top}x_{t}\right]^{-1}_{jj}}$ (30)
To estimate the standard error $SE(\hat{\beta})$, the time-level population variance is needed $\sigma^{2}_{t}$ (Domínguez Islas & Rice, 2018). Grouping by time, $t$ instead of by study, $s$, is a key difference since variation in time is now explicitly controlled for, instead of being implicitly controlled for by the grouping of studies. The subscript $jj$ refers to the $j-th$ diagonal elements of the variance-covariance matrix of the estimator, this syntax is used when there are multiple covariates.
Combined Location and Time-level Fixed Effects
The estimation of parameters and their standard errors will be calculated similarly as in Section 4.1 & 4.2 with the same estimation strategy. Both indices need to be paired together since each observation is observed at a specific location and time period. The parameter is estimated using the weight of both groupings (location and time) as seen in Equation 31 below.
Appendix H - Details on MSE & Similar Measures
As mentioned in Section 5.2.2, a common metric the meta-analysis simulation literature measures bias is with mean squared error (MSE) $\frac{1}{n}\sum^{N}_{i=1}(\hat{\beta}_{i}-\beta_{i})^{2}$. But, there are other variations which are rare in this literature, but well circulated in statistics and econometrics literature. These include mean absolute error, mean percent error, or mean absolute percent error (respectively MAE, MPE or MAPE) (Gneiting, 2011; West, 2006; Timmermann, 2006; Maronna et al., 2019; Hyndman & Athanasopoulos, 2021). Note that estimator bias is used instead of these variations because our simulation uses all the in-sample data to estimate, so these would just calculate model performance instead of the intended bias. While MSE is designed to penalize larger deviations from the true value, MAE is designed without penalty via $\frac{1}{n}\sum^{N}_{i=1}|\hat{\beta}_{i}-\beta_{i}|$. Section 5.1 shows that our simulation is not designed to have outliers, but if it did, MAE would be more appropriate to use than MSE due to the sensitivity of MSE to outliers (Maronna et al., 2019; Müller & Guido, 2017; Hyndman & Athanasopoulos, 2021; French et al., 1987). MSE is scale dependent, so if this is a concern it may be more appropriate to use MPE which is scale independent and measures the percentage difference via $\frac{1}{n}\sum^{N}_{i=1}\frac{\hat{\beta}_{i}-\beta_{i}}{\beta_{i}}$ (Gneiting, 2011; West, 2006; Diebold & Lopez, 1996). MAPE is simply the absolute value of the percentage difference from MPE which in addition to being scale independent it focuses on the magnitude of prediction error instead of over/under prediction like MPE (Gneiting, 2011; West, 2006; Timmermann, 2006).